# WaterSeeker: Efficient Detection of Watermarked Segments in Large Documents

### v2

@@ 5.2 Suspicious Region Localization
  1. 1.
    Score List Computation: We utilize a small sliding window (i.e., $W=50$) to efficiently traverse the entire text. For each position of the window, we compute a statistical measure that reflects the average watermark intensity. For example, in the KGW method, this statistic is represented as $|s|_{G}/W$, indicating the density of green tokens inside the window. Meanwhile, in the Aar method, it is calculated as $\cfrac{\sum\log(1/(1-u))}{W}$, representing the average correlation value between tokens and the pseudo-random sequence within the window. This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity of $n_{i}$ to $n_{i}+W$.
  2. 2.
    Anomaly Extraction: In this step, we design an anomaly extraction algorithm to identify outliers from the score list. We calculate the mean of the score list, the top-k mean, and the standard deviation, denoted as $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$, respectively. The extracted outliers are the points that satisfy the following condition, where $\theta_{1}$ and $\theta_{2}$ are hyperparameters:
    $\text{score}>s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2}).$ (7)
@@
    Fragment Connection: In this step, we use a connecting method with a certain tolerance to link nearby outliers, then filter out segments that are too short, returning a list of indices.
The scores within watermarked segments tend to be relatively stable and close to the top-k mean. Concurrently, the scores corresponding to segments outside the watermark also remain stable and approximately equal to the overall mean (because a large volume of natural text dilutes the overall watermark strength). As a result of these characteristics, when we use $(s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1}\quad(\theta_{1}<1)$ as a criterion for distinction, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy the following conditions: $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$. This ensures that the extracted suspicious watermarked regions are likely to cover the actual segments while keeping the starting and ending deviations within a window size.
Algorithm 1 WaterSeeker Algorithm
2:     Compute score list using sliding window
3:     Calculate $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$
4:     Detect anomalies using threshold:
6:     Connect nearby outliers and filter short segments
7:     return filteredSegments
8:end procedure
10:     hasWatermark $\leftarrow$ False
11:     indices $\leftarrow$ [ ]
12:     for $(s^{\prime},e^{\prime})$ in suspiciousRegions do
13:         maxScore = $\infty$- ∞, maxIndice = ()
15:              if DetectWatermark(tokens[$s:e$])$>$ maxScore then
16:                  maxScore = DetectWatermark(tokens[$s:e$])
17:                  maxIndice = $(s,e)$
18:              end if
19:         end for
20:         if maxScore $>$ threshold then
21:              hasWatermark $\leftarrow$ True
22:              indices.append(maxIndice)
23:         end if
24:     end for
25:     return hasWatermark, indices
26:end procedure
30:     return hasWatermark, indices
31:end procedure
5.3 Local Traverse Detection
After obtaining the localization results from the first step, the second phase (fine step), local traverse detection, conducts a more detailed verification. For each $(s^{\prime},e^{\prime})$ pair in the localization results, traverse inwards through the segments where the start point falls within $[s^{\prime},s^{\prime}+W)$ and the end point within $(e^{\prime}-W,e^{\prime}]$. Based on the previous analysis of the ranges of $s^{\prime}$ and $e^{\prime}$, this traversal has a high probability of reaching the gold index. Perform full-text detection sequentially on these segments, and select the statistic with the most significance (e.g., for KGW, choose the segment with the largest z-score, and for Aar, the segment with the smallest p-value) to compare with the threshold. The pseudocode for the entire WaterSeeker algorithm can be found in Algorithm 1.
5.4 Time Complexity Analysis
Time Complexity Analysis of WaterSeeker:
  1. 1.
    Suspicious Region Localization: $O(N)$. This step involves three sub-steps: score list computation, anomaly detection, and fragment connection. Each of these sub-steps has a time complexity of $O(N)$, where $N$ is the length of the text.
  2. 2.
    Local Traverse Detection: $O(W^{2})$. For each suspicious region identified, this step examines $W^{2}$ windows, where $W$ is the window size.
The total time complexity of WaterSeeker is $O(N+W^{2})$. In practice, $W^{2}$ is typically maintained at a value lower than $T$, as even a slightly larger window (i.e., $W=50$) is sufficient to provide a relatively smooth and low-noise representation of the surrounding watermark intensity. Therefore, the dominant term in most cases is $O(N)$, making the overall time complexity of WaterSeeker $O(N)$.
Lower Bound Complexity Analysis for the Problem: To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once. This necessitates at least one full pass through the text, establishing a lower bound of $\Omega(N)$ for the time complexity of this problem.
In conclusion, the WaterSeeker algorithm, with its $O(N)$ complexity in typical cases, achieves a time complexity that matches the theoretical lower bound of the problem, indicating that it is asymptotically optimal for the task of detecting watermarked segments in long texts.
6 Experiment
6.1 Experiment Settings
Watermarking Methods and Language Models: We selected two representative watermarking algorithms, KGW (Kirchenbauer et al., 2023) and Aar (Aaronson & Kirchner, 2022), each with three strength levels. For KGW, the strength is determined by parameter $\delta$, with values 2.0, 1.5, and 1.0 (from strong to weak). For Aar, the strength is controlled by parameter temperature, with values 0.3, 0.2, and 0.1 (from strong to weak). We used Llama-2-7b (Touvron et al., 2023) and Mistral-7b (Jiang et al., 2024) as generation models.
Dataset Construction: Following previous studies, the C4 dataset (Raffel et al., 2020) was used for prompts, with the first 30 tokens of each entry serving as input. Watermarked segments of random length (100 to 400 tokens) were generated using randomly selected watermark strengths. For positive examples, one such segment was randomly inserted into each 10,000-token Wikipedia passage (Foundation, ). Negative examples consist of unmodified 10,000-token Wikipedia corpus. Based on this procedure, four datasets were created, each containing 300 positive and 300 negative examples: KGW-llama, KGW-mistral, Aar-llama, and Aar-mistral. These datasets employ different combinations of watermarking methods and LLMs. The specific distributions of watermark strengths and lengths are detailed in Appendix D.
It is important to note that watermarked segments with higher strength and longer length are inherently easier to detect and demonstrate better robustness against the dilution effect. Therefore, to showcase the superiority and adaptability of WaterSeeker, each dataset incorporates watermarked segments of varying strengths and lengths.
Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100. The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms. The threshold selection within the specified window is detailed in Appendix C. Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives. Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
6.2 Detection Performance Analysis
@@
Table 1: We compared the detection performance of WaterSeeker with other methods, including Plain Detect, WinMax (Kirchenbauer et al., 2024), and FLSW. The reported metrics include False Positive Rate (FPR), False Negative Rate (FNR), F1 score, and the average Intersection over Union (IoU) between detected segments and ground truth segments in positive samples.
@@
FPR FNR F1 IoU FPR FNR F1 IoU
@@
WaterSeeker(Ours) 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets. Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection. Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red.
The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax. Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance. However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4). Furthermore, the IoU [^1] For samples with no detected segments, the IoU is 0. For samples with multiple detected segments, take the max(IoU). results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
6.3 Adaptability Analysis
While the four datasets used in the main experiment incorporated a mix of watermarked segments with varying strengths and lengths, this section presents a more granular comparison of the detection capabilities of WaterSeeker and the FLSW method (FLSW-100, FLSW-200, FLSW-300, and FLSW-400) across specific watermark strengths and lengths. As shown in Figure 3, the performance of different detection algorithms exhibits a consistent trend: watermarked segments with higher strengths and longer lengths are more easily detected. Moreover, it is evident that FLSW performs well within a length range close to its window size, but shows poor adaptability in other ranges. This limitation arises from its fixed-length nature, which leads to dilution effects when detecting longer or shorter watermarked segments, thereby reducing detection performance. In contrast, WaterSeeker exhibits good adaptability across different length ranges by using anomaly extraction techniques. Additionally, it incorporates a top-k score mechanism to adapt to varying watermark strengths.
@@ 6.3 Adaptability Analysis
6.4 Time Complexity Analysis
Table 2: Average detection time per sample for various detection methods ($T\approx 10,000$, Unit: s).
@@ 6.4 Time Complexity Analysis
Table 2 lists the average time spent per sample for WaterSeeker and other baseline algorithms during detection. The results show that the time taken by WaterSeeker is comparable to that of full-text detection and FLSW, and significantly lower than the time required by WinMax. Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively. This section aims to offer a more comprehensive analysis based on experimental results, taking into account the constant factors preceding the complexity terms. Additionally, it will investigate the reasons behind the significant disparity in execution time of WinMax between KGW and Aar.
In this experimental setting, let the time required to compute a score in the score list be $t_{1}$, and the time required to compute the watermark detection statistic within a local window be $t_{2}$. The time expenditure for the WinMax [^2] The WinMax algorithm, in practical implementation, also requires computing the score list for the entire text first, then traversing the windows to calculate the detection statistic. algorithm to detect a sample can be expressed as $N\cdot t_{1}+N\cdot(W_{max}-W_{min})\cdot t_{2}$. For WaterSeeker, the first step of localization consists of three sub-steps: score list computation, anomaly detection, and fragment connection. Each of these sub-steps has a complexity of $O(N)$. If we denote the time required for the latter two sub-steps as $t_{3}\cdot N$, then the total time for the first step is $N\cdot t_{1}+N\cdot t_{3}$. The second step, local traverse detection, requires $W^{2}\cdot t_{2}$ time. Consequently, the total time expenditure for the entire process can be expressed as $N\cdot t_{1}+N\cdot t_{3}+W^{2}\cdot t_{2}$.
For the Aar (Aaronson & Kirchner, 2022) watermarking method, calculating the watermark detection statistic involves computing a GammaTransform function, which is a relatively complex computation, resulting in a larger value for $t_{2}$. In this case, the time difference between WinMax and WaterSeeker will become significant due to the large disparity in the coefficients preceding $t_{2}$. In this experiment, with $T=10,000,W_{max}=400,W_{min}=100,W^{2}=2,500$, the coefficient difference amounts to a factor of 1,200. For the KGW (Kirchenbauer et al., 2023) watermarking method, calculating the z-score within a local window is a relatively straightforward computation. In this case, the time difference between WinMax and WaterSeeker will be reduced.
Through the above analysis, it is clear that WaterSeeker is a general watermark detection method with consistently low time expenditure across various watermarking algorithms. In contrast, WinMax exhibits extremely high time costs in situations where calculating the watermark detection statistic is complex, making it challenging for practical use.
6.5 Ablation Study
@@
In this section, we will analyze the effectiveness of the two stages of WaterSeeker through an ablation study. The first stage of WaterSeeker, Suspicious Segment Localization, aims to achieve significant coverage of the ground truth segments while keeping the start and end offsets within a specified window size. This ensures that subsequent local traversals can reach the gold index. Table 3 illustrates the performance of Step 1, showing that the localization results of Suspicious Segment Localization achieve an average coverage level exceeding 0.95 across various watermark algorithms and strengths. Additionally, both the average start offset and average end offset are less than 50, remaining within the designated window size. This indicates that Step 1 fulfills its intended purpose effectively.
Local Traverse Detection performs a localized traversal based on the segments narrowed down in the first step, allowing for more refined verification within the window. Table 4 illustrates the differences in detection outcomes when using Local Traverse compared to directly applying full-text detection with the localization results. It can be found that across different LLMs and watermarking algorithms, Local Traverse consistently enhances the detection performance, making it an indispensable component of WaterSeeker.
Table 4: This table compares the detection performance of using Local Traverse versus not using it across the four datasets.
@@
FPR FNR F1 IoU FPR FNR F1 IoU
Llama-2-7b w. local traverse 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
@@
Mistral-7b w. local traverse 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
@@
7 Further Application

### v3

@@ 5.2 Suspicious Region Localization
(1) Score List Computation: A small sliding window (i.e., $W=50$) is used to traverse the entire text. For each window position, a statistical measure is computed to reflect the average watermark intensity. In the KGW method, this statistic is $|s|_{G}/W$, representing the density of green tokens within the window. In the Aar method, it is calculated as $\frac{\sum\log(1/(1-u))}{W}$, representing the average correlation value between tokens and the pseudo-random sequence within the window. This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity from $n_{i}$ to $n_{i}+W$.
(2) Anomaly Extraction: In this step, we design an anomaly extraction algorithm to identify outliers from the score list. We calculate the mean of the score list, the top-k mean, and the standard deviation, denoted as $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$, respectively. The extracted outliers are the points that satisfy the following condition, where $\theta_{1}$ and $\theta_{2}$ are hyperparameters:
$\text{score}>s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2}).$
@@
(3) Fragment Connection: In this step, we use a connecting method with a certain tolerance to link nearby outliers, then filter out segments that are too short, returning a list of indices.
@@
Table 1: We compared the detection performance of WaterSeeker with other methods, including Full-text Detection, WinMax (Kirchenbauer et al., 2024), and FLSW. The reported metrics include FPR, FNR, F1 score, and the average IoU between detected segments and ground truth segments in positive samples.
@@
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
@@
WaterSeeker(Ours) 0.007 0.290 0.827 0.641 0.010 0.390 0.753 0.542
The scores within watermarked segments are relatively stable and close to the top-k mean, while scores outside the watermark remain stable and close to the overall mean due to the dilution effect of large natural text. When using $(s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1}\quad(\theta_{1}<1)$ as a distinction criterion, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$. This ensures that the extracted suspicious watermarked regions likely cover the actual segments, with starting and ending deviations within a window size.
5.3 Local Traverse Detection
After obtaining the localization results from the first step, the second phase (fine step), local traverse detection, conducts a more detailed verification. For each $(s^{\prime},e^{\prime})$ pair in the localization results, traverse inwards through the segments where the start point falls within $[s^{\prime},s^{\prime}+W)$ and the end point within $(e^{\prime}-W,e^{\prime}]$. Based on the previous analysis of the ranges of $s^{\prime}$ and $e^{\prime}$, this traversal has a high probability of reaching the gold index. Perform full-text detection sequentially on these segments, and select the statistic with the most significance to compare with the threshold. The pseudocode for the entire WaterSeeker algorithm can be found in Algorithm 3.
5.4 Time Complexity Analysis
Time Complexity of WaterSeeker: Suspicious Region Localization has a time complexity of $O(N)$. This step involves score list computation, anomaly detection, and fragment connection, each with a time complexity of $O(N)$, where $N$ is the text length. Local Traverse Detection has a time complexity of $O(W^{2})$. For each suspicious region, this step examines $W^{2}$ windows, where $W$ is the window size. The total time complexity of WaterSeeker is $O(N+W^{2})$. In practice, $W^{2}$ is typically kept lower than $N$, as a slightly larger window (i.e., $W=50$, detailed in Appendix E) suffices for a smooth and low-noise representation of the surrounding watermark intensity. Thus, the overall time complexity of WaterSeeker is $O(N)$.
Lower Bound Complexity for the Problem: To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once. This requirement establishes a lower bound of $\Omega(N)$ for the time complexity of the problem, as at least one full pass through the text is necessary. Consequently, the WaterSeeker algorithm achieves a time complexity that matches the theoretical lower bound of the problem.
6 Experiment
@@
Table 2: Average detection time per sample for various detection methods ($N\approx 10,000$, Unit: s).
@@
6.1 Experiment Settings
Watermarking Methods and Language Models: We selected two representative watermarking algorithms, KGW (Kirchenbauer et al., 2023) and Aar (Aaronson and Kirchner, 2022), each at three strength levels. KGW’s strength was set by the $\delta$ parameter (2.0=strong, 1.5=medium, 1.0=weak), while Aar’s strength used the temperature parameter (0.3=strong, 0.2=medium, 0.1=weak). We used Llama-2-7b (Touvron et al., 2023) and Mistral-7b (Jiang et al., 2024) as generation models.
Dataset Construction: The first 30 tokens of each entry in the C4 dataset (Raffel et al., 2020) were used for prompts. Watermarked segments of random length (100 to 400 tokens) were then generated using randomly selected watermark strengths. For positive examples, one such segment was randomly inserted into each 10,000-token Wikipedia passage (Foundation, ). Negative examples consist of unmodified 10,000-token Wikipedia corpus. Based on this procedure, four datasets were created, each containing 300 positive and 300 negative examples: KGW-llama, KGW-mistral, Aar-llama, and Aar-mistral. The specific distributions of watermark strengths and lengths are detailed in Appendix F.
Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100. The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms. The threshold selection within the specified window is detailed in Appendix D. Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives. Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
@@ 6.1 Experiment Settings
Table 4: Comparison of detection performance with and without Local Traverse Detection across four datasets.
@@
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b w. local traverse 0.017 0.313 0.806 0.624 0.010 0.440 0.713 0.507
@@
Mistral-7b w. local traverse 0.007 0.290 0.827 0.641 0.010 0.390 0.753 0.542
@@
6.2 Detection Performance Analysis
In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets. Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection. Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red. Additional performance results for WaterSeeker under various mixing ratios are presented in Appendix G.
The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax. Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance. However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4). Furthermore, the IoU results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
6.3 Adaptability Analysis
While the main experiment used four datasets with a mix of watermarked segments of varying strengths and lengths, this section compares the detection capabilities of WaterSeeker and the FLSW method across specific watermark strengths and lengths. Figure 3 shows that the performance of different detection algorithms follows a consistent trend: watermarked segments with higher strengths and longer lengths are more easily detected. FLSW performs well within a length range close to its window size but shows poor adaptability in other ranges due to its fixed-length nature, which leads to dilution effects when detecting longer or shorter watermarked segments. In contrast, WaterSeeker exhibits good adaptability across different length ranges by using anomaly extraction techniques and incorporates a top-k score mechanism to adapt to varying watermark strengths.
6.4 Time Complexity Analysis
Table 2 shows the average time spent per sample for WaterSeeker and other baseline algorithms during detection. WaterSeeker’s time is comparable to full-text detection and FLSW, and significantly lower than WinMax. Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively. This section offers a more comprehensive analysis based on experimental results, considering constant factors and investigating the significant disparity in WinMax’s execution time between KGW and Aar.
Let $t_{1}$ be the time to compute a score in the score list, and $t_{2}$ be the time to compute the watermark detection statistic within a local window. WinMax’s time expenditure to detect a sample is $N\cdot t_{1}+N\cdot(W_{\max}-W_{\min})\cdot t_{2}$. For WaterSeeker, the localization step consists of score list computation, anomaly detection, and fragment connection, each with complexity $O(N)$. Denoting the time for the latter two sub-steps as $t_{3}\cdot N$, the total time for the first step is $N\cdot t_{1}+N\cdot t_{3}$. The local traverse detection step requires $W^{2}\cdot t_{2}$ time. Thus, WaterSeeker’s total time expenditure is $N\cdot t_{1}+N\cdot t_{3}+W^{2}\cdot t_{2}$.
For the Aar (Aaronson and Kirchner, 2022) watermarking method, calculating the watermark detection statistic involves a complex GammaTransform function, resulting in a larger $t_{2}$. This leads to a significant time difference between WinMax and WaterSeeker due to the large disparity in coefficients preceding $t_{2}$. In this experiment, with $N=10,000,W_{\max}=400,W_{\min}=100,W^{2}=2,500$, the coefficient difference is a factor of 1,200. For the KGW (Kirchenbauer et al., 2023) watermarking method, calculating the z-score within a local window is relatively straightforward, reducing the time difference between WinMax and WaterSeeker.
WaterSeeker is a general watermark detection method with consistently low time expenditure across various watermarking algorithms. In contrast, WinMax exhibits extremely high time costs when calculating the watermark detection statistic is complex, making it challenging for practical use.
6.5 Ablation Study
We analyze the effectiveness of the two stages of WaterSeeker through an ablation study. The first stage, Suspicious Segment Localization, aims to achieve high coverage of the ground truth segments while keeping the start and end offsets within a specified window size. This ensures subsequent local traversals can reach the gold index. Table 3 shows that Step 1 achieves an average coverage exceeding 0.95 across various watermark algorithms and strengths, with average start and end offsets less than 50, remaining within the designated window size. Step 1 fulfills its purpose effectively.
Local Traverse Detection performs a localized traversal based on the segments from Step 1, allowing for more refined verification within the window. Table 4 shows that across different LLMs and watermarking algorithms, Local Traverse consistently enhances detection F1 score and average IoU compared to directly applying full-text detection with the localization results, making it an indispensable component of WaterSeeker.